My brother is 23 years old and my sister is 20 years old. Find the sum of their ages when the ratio of their ages is 3:2?
Answer by Jk22(389)

when the ratio is 3:2, hence with an offset of n years, then (23+n)/(20+n)=3/2

we have then to find 23+n+20+n=43+2n

we need to find n from the first equation : 2*(23+n) = 3(20+n)

46+2n = 60 + 3n, or n = 46-60 = -14 years. This was the case 14 years ago.

then the sum is : 6+9= 15 years.